OpenStudy (study_buddy99):

Algebra II: Will medal- solving rational equations (question in description)

OpenStudy (study_buddy99):

\[\frac{ x^2-3x-4 }{x^3-x^2 }-\frac{ 1 }{ x^2 }=\frac{ x-2 }{ x^2 }\]

OpenStudy (danjs):

remember the other question, get rid of those fractions first, multiply everyting by a common denominator, i would factor the x^3-x^2 first to x^2(x-1) ... maybe?

OpenStudy (study_buddy99):

I'm stuck at this part though:

OpenStudy (faiqraees):

Take x^2 common in the denominator of the first fraction on left hand side

OpenStudy (faiqraees):

or you can start by taking 1/x^2 on the right hand side of the equation and then simplifying the fractions on right hand side

OpenStudy (danjs):

\[\large\frac{ x^2-3x-4 }{x^2(x-1) }-\frac{ 1 }{ x^2 }=\frac{ x-2 }{ x^2 }\] multiply everything by x^2(x-1) now

OpenStudy (faiqraees):

no that will make it complicated

OpenStudy (faiqraees):

\[\large\frac{ x^2-3x-4 }{x^2(x-1) }-\frac{ 1 }{ x^2 }=\frac{ x-2 }{ x^2 } \]\[\large\frac{ x^2-3x-4 }{x^2(x-1) }=\frac{ x-2 }{ x^2 }+\frac{ 1 }{ x^2 } \]\[\large\frac{ x^2-3x-4 }{x^2(x-1) }=\frac{ x-2+1}{ x^2 } \]

OpenStudy (faiqraees):

Start like this

OpenStudy (study_buddy99):

\[\frac{ x^2-3x-4(x^3-x^2) }{x^3-x^2 }-\frac{ -1(x^3-x^2) }{ 1x^2 }=\frac{ x-2(x^3-x^2) }{ x^2 }\]

OpenStudy (danjs):

whichever way you want, try both

OpenStudy (faiqraees):

\[\large\frac{ x^2-3x-4 }{x^2(x-1) }=\frac{ x-1 }{ x^2 } \] Now multiply x^2 on both sides \[\large\frac{ x^2-3x-4 }{(x-1) }=x-1 \] Here you have completely reduced fraction

OpenStudy (study_buddy99):

hm...

OpenStudy (study_buddy99):

but I mean, how does one cancel the negative and positive relationship shown on the second fraction?

OpenStudy (danjs):

\[\large\frac{ x^2-3x-4 }{x^2(x-1) }-\frac{ 1 }{ x^2 }=\frac{ x-2 }{ x^2 }\] multiply everything bythe greatest common denominator x^2(x-1) \[(x^2-3x-4) - 1(x-1)=(x-2)(x-1)\]

OpenStudy (study_buddy99):

okay, I see

OpenStudy (danjs):

can you expand all that out , distribute everything

OpenStudy (danjs):

remember the end goal is just to get , x , alone and equal someting... x...

OpenStudy (study_buddy99):

I think I've almost got it

OpenStudy (danjs):

\[\large x^2-3x-4 - x + 1 = x^2-x-2x+2\]

OpenStudy (danjs):

easier, the x^2 cancel away

OpenStudy (danjs):

type what you doin, not sure

OpenStudy (danjs):

hmm..here is the dirstibution to expand \[(x^2-3x-4) - 1(x-1)=(x-2)(x-1)\] 

OpenStudy (study_buddy99):

oh I see, so my answer would be -5?

OpenStudy (danjs):

i typed that wrong with the + - sines x^2 - 3x - 4 - x + 1 = x^2 - 3x + 2

OpenStudy (danjs):

take an x^2 off both sides, it cancels out, good combine all the like x terms and the constant number terms -4x - 3 = -3x+2

OpenStudy (danjs):

add 4x both sides -3 = x + 2 take off 2 -5 = x yep

OpenStudy (study_buddy99):

open study is glitching bad tonight and making me repeat previous conversation. Anyway, thank you so much!

OpenStudy (danjs):

welcome, just moving things around in an equation... a few practice probs and you have it down